Mr. Thomas is putting up a fence around his field. The field is 360yd long and 225yd wide. He wants to put up a fence post every 3yd. How many fence posts will me. Thomas need?
Answer by Theo(13342)

the 360 yard long piece will require 360 / 3 = 120 + 1 = 121 posts.
the 225 years piece will require 225 / 3 = 75 posts - 1 = 74 posts.
you get a total of 195 posts.
double that and you get 390.

consider it was a very small field that was 9 yards long by 6 yards wide.
you would require 9/3 = 3 + 1 = 4 posts for the length
you would require 6/3 = 2 - 1 = 1 for the width.
sounds counter intuitive but this is what it will look like.

                  P     P     P     P

               
                  P                 P


                  P     P     P     P

the corner posts are already placed in when you put the posts in for the length.
that means that you don't need posts on the corner for the width because they are already there when you put them in for the length.
in between each P is a 3 foot section.
there are three 3 foot sections on the length.
there are two 3 foot sections on the width.